OpenStudy (anonymous):

What is the lateral area of a triangular prism with base edges all equaling 3 meters and a height of 5 meters?

OpenStudy (anonymous):

at first the base area s1=(1/2)*a*b*sinr=1/2*3^2*sin60=\[(9\sqrt{3})/4\] and the side area s2=5*3*3=45 s=s1+s2=45+\[(9\sqrt{3})/4\]